# Path integral and conformal instability in nonlocal quantum gravity

### v1

@@ 3 Tree-level graviton propagator
can be computed in a standard albeit lengthy way. The expansion at second order of the action (4) yields the Lagrangian [^2] Here is how to compare the conventions and results of Accioly:2002tz with ours. Spacetime signature is mostly minus in Accioly:2002tz and mostly plus for us. Also, their Ricci scalar $R$ is $-R$ for us, so that there is an extra overall -- sign in the total action and in its coefficients $\alpha=-M_{\text{\tiny Pl}}^{2}\gamma_{0}$, $\beta=-M_{\text{\tiny Pl}}^{2}\gamma_{2}$ and $\gamma=-M_{\text{\tiny Pl}}^{2}\gamma_{4}$. Then, $b=\beta\kappa^{2}/2=-2\gamma_{2}$, $c=\alpha/\beta=\gamma_{0}/\gamma_{2}$ and $d=\gamma\kappa^{2}/2=-2\gamma_{4}$, where $\kappa=2/M_{\text{\tiny Pl}}$. Finally, the graviton in Accioly:2002tz is dimensionful and defined as $g_{\mu\nu}=\eta_{\mu\nu}+\kappa h_{\mu\nu}$, while our $h_{\mu\nu}$ is dimensionless. Therefore, our coefficients $c_{i}\in\{c_{1},c_{2},c_{0},\bar{c}_{0},\bar{\bar{c}}_{0}\}$ are related to the coefficients $x_{i}\in\{x_{1},x_{2},x_{0},\bar{x}_{0},\bar{\bar{x}}_{0}\}$ in Accioly:2002tz by $c_{i}(k^{2})=-(M_{\text{\tiny Pl}}^{2}/4)\,x_{i}(k^{2})$. The expressions in Accioly:2002tz immediately generalize to the weakly nonlocal case with nontrivial form factors and we have recalculated them anew.
@@
Contrary to $\mathcal{O}^{\rm K}$, the kinetic operator $\mathcal{O}=\mathcal{O}^{\rm K}+\mathcal{O}^{\rm gf}$ is invertible. In the basis of Barnes–Rivers projectors VanNieuwenhuizen:1973fi and in momentum space, it is
@@
This expression is equivalent ${\mathcal{O}}_{\rm K}^{-1}={P}^{(2)}/c_{2}+{P}^{(0)}/c_{0}$ when setting $\lambda_{1}=\lambda_{2}=\lambda_{3}=0$. In general, different gauge choices become handy when exploring different aspects of the theory, for instance, when showing that the spin-0 mode with the wrong sign does not propagate or positivity of the eigenvalues of the kinetic operator in Euclidean signature. Some common gauge choices are the Julve–Tonin gauge ($\lambda=0$) Julve:1978xn , the de Donder gauge ($\lambda=1/2$, $\lambda_{2}=0=\lambda_{3}$), and the Feynman gauge ($\lambda=1/2$, $\lambda_{1}=1$, $\lambda_{2}=0=\lambda_{3}$) Accioly:2002tz .
@@ 4 Conformal instability
where the fields $\Phi_{i}$ include the metric $g_{\mu\nu}$ and $S_{\rm E}$ is the Euclidean action. This theory can suffer from a conformal instability problem Hawking:1976ja ; Gibbons:1978ac ; Ham09 , which we illustrate in Einstein gravity. A conformal transformation of the metric
@@
in the Ricci term $R$ produces a kinetic term for the conformal factor $\Omega$ with the wrong sign:
$\frac{M_{\text{\tiny Pl}}^{2}}{2}\int d^{D}x\,\sqrt{|g|}\,R=\frac{M_{\text{\tiny Pl}}^{2}}{2}\int d^{D}x\,\sqrt{|\hat{g}|}\,\Omega^{D-4}\left[\Omega^{2}\hat{R}+(D-1)(D-2)\hat{\partial}_{\mu}\Omega\hat{\partial}^{\mu}\Omega\right],$ (69)
@@
Therefore, the Euclidean path integrand $\exp(iS)=\exp(-S_{\rm E})$ has an anti-Gaussian term $\sim\exp[\int(\partial\omega)^{2}]$ and is unbounded from above. To avoid this arbitrarily large instability, one can deform the integration contour over the fields, in particular, taking a functional measure $[\mathcal{D}g_{\mu\nu}]$ over complex Euclidean metrics (complex conformal factor), in such a way as to get a convergent result Gibbons:1978ac .
@@
The physical significance of this procedure was criticized in Schleich:1987fm ; Hartle:2020glw ; Mazur:1989by and motivated another approach, where the fundamental definition of the path integral is Lorentzian and one derives (instead of assuming) the correct Euclidean path integral by integrating over a suitable contour in field space with real Lorentzian metrics Schleich:1987fm ; Hartle:2020glw ; Mazur:1989by ; Marolf:2022ybi . Nonlocal quantum gravity is defined with a Lorentzian path integral, so that the issue of the conformal instability can be solved in this way. At least, there is no reason to suspect that this resolution, successful in other theories, should not work also here.
@@
In this paper, we will not discuss how such a contour is deformed because we will mainly focus on perturbation theory, where the Lorentzian path integral is employed. However, amplitudes are calculated in Euclidean momenta, as we discussed above. Obviously, in Euclidean quantum gravity, the first step 1. in section 2.5 would be automatically implemented but the final result would be distinguishable from what obtained in Lorentzian theory because correlation functions (hence scattering amplitudes) are calculated from the path integral, and the latter (together with its functional measure) differs in the Euclidean and Lorentzian formulations. Redefining and analytically continuing the Lorentzian momenta of the fields appearing in the functional measure of the Lorentzian path integral produces a measure which is functionally inequivalent to the one obtained by analytically continuing the Euclidean momenta of the fields appearing in the measure of the Euclidean path integral.
@@
From this discussion, it stems that the conformal instability problem does not affect perturbative results because, in general, the Lorentzian perturbation theory is well defined Ham09 . Let us expand on this statement. In order for the Lorentzian path integral to be well defined in perturbation theory, the tree-level graviton kinetic term in Feynman prescription $k^{2}\to k^{2}-i\epsilon$ must generate a Gaussian term in the path integral that makes it convergent. If this was the case, then we would have convergence order by order in perturbation theory because, at higher orders, one only has to include vertex insertions, which do not spoil the tree-level convergences property. This is because all correlation functions in perturbation theory are written as an expansion of the interactions $S_{\rm int}$: schematically for a generic field $\Phi$,
@@
where $S_{0}=\int\Phi\mathcal{O}\Phi$ and $\mathcal{O}=\mathcal{O}^{\rm K}+\mathcal{O}^{\rm gf}$ is the tree-level kinetic term (inverse of the propagator), where $\mathcal{O}^{\rm K}$ is the kinetic term coming from the bare action and $\mathcal{O}^{\rm gf}$ is a contribution coming from the gauge-fixing action. In order to explicitly compute any perturbative amplitude, we should commute the functional integral with the sum. This issue is related to the Borel summability of the perturbative series that deserves to be investigated also in gravity and, in particular, in a finite theory of quantum gravity in the quantum field theory framework.
If one naively considered only the gauge-invariant part of the propagator $\mathcal{O}^{-1}_{\rm K}$, the convergence problem would persist after the $k^{2}\to k^{2}-i\epsilon$ prescription because the spin-2 and spin-0 modes in $\mathcal{O}^{\rm K}$ have opposite sign, so that the spin-0 mode would generate an anti-Gaussian profile. Thus, the role of $\mathcal{O}^{\rm gf}$ cannot be ignored. If, after a Weyl transformation (68), a suitable gauge choice exists such that the eigenvalues of the kinetic operator of the conformal factor $\Omega$ are all zero or negative in Euclidean signature, then there is convergence of the Euclidean path integral at the perturbative level.
To show this, it is sufficient to consider the gravitational sector of nonlocal quantum gravity and the generic action (4) where now $\gamma_{0}$, $\gamma_{2}$ and $\gamma_{4}$ are generic form factors which vanish for Einstein gravity and are constant in Stelle gravity, and $S_{\rm gf}$ is the gauge-fixing action. In a conformally invariant theory, the metric $g_{\mu\nu}$ composing the measure and the curvature tensors is the one before making explicit the dilaton dependence. Decompose the metric as in (57) and derive the perturbed action for the dimensionless spin-2 field $h_{\mu\nu}$ (the graviton polarization modes are inside this object). On Minkowski background, the kinetic term is $\propto h^{\mu\nu}\mathcal{O}_{\mu\nu\sigma\tau}h^{\sigma\tau}$, where the operator $\mathcal{O}_{\mu\nu\sigma\tau}$ is given in section 3 and is a generalization of the higher-derivative expression found in Accioly:2002tz .
@@
where we used eq. (60), the Barnes–Rivers projectors (3), and the ensuing properties
@@
Intuitively, only the spin-0 projectors survive because they act on a scalar. From the expressions (3) for the coefficients $c_{0}$, $\bar{c}_{0}$, and $\bar{\bar{c}}_{0}$, in momentum space we have
@@
If we ignored the gauge-fixing terms and set $\lambda=\lambda_{1}=\lambda_{2}=0$, then $\mathcal{K}>0$ in Euclidean signature for non-negative form factors $\gamma_{0,2}\geqslant 0$ and the Euclidean action $S_{\rm E}\sim-\int\phi\mathcal{K}\phi$ would be unbounded from above. This is the conformal instability problem. However, by virtue of gauge invariance, if we can find a range of gauge choices where $\mathcal{K}\leqslant 0$ for all momenta $k$, then we can conclude that such problem is a gauge artefact and that the scalar mode $\phi$ does not propagate ($\mathcal{K}=0$) or that it propagates with a kinetic term of the “straight” sign ($\mathcal{K}<0$). In general, for $\gamma_{2}^{\prime}\neq 0$ this happens when
@@
The first of these conditions appears in all theories with an Einstein–Hilbert term and implies that both $\lambda$ and $\lambda_{1}$ can be chosen to be constant. In $D=4$ dimensions, $\lambda_{1}\geqslant 3/(4\lambda-1)^{2}$ and any gauge with $\lambda\neq 1/4$ can fulfill eq. (76), including Julve–Tonin ($\lambda=0$) and de Donder ($\lambda=1/2$, $\lambda_{2}=0=\lambda_{3}$), but not the Feynman gauge ($\lambda=1/2$, $\lambda_{1}=1$, $\lambda_{2}=0=\lambda_{3}$).
@@ 5 Discussion
To date, nonlocal quantum gravity has been studied only with perturbative techniques in relation with scattering amplitudes, perturbative unitarity, renormalization, black-hole solutions and cosmology. However, the conformal instability problem could arise when considering nonperturbative processes and the full path integral not expanded in terms of the interactions. As in the perturbative case, once the Euclidean path integral is derived from the Lorentzian one, a nontrivial choice of the gravitational measure of the path integral can make it convergent. The physical interpretation is that this measure would enhance the weight of strong metric fluctuations pushing away physical fluctuations from the conformal unbounded abyss Ham09 . This seems indeed to be the case, both on a lattice at strong coupling Hamber:1984kx ; Berg:1985ka and from a nonperturbative extension Dasgupta:2001ue ; Ambjorn:2002gr of the perturbative calculation of Mazur:1989by . In the case of the Einstein–Hilbert action, after a Wick rotation of the nonperturbative Lorentzian path integral defined on a causal dynamical triangulation, it turns out that the term responsible for the conformal instability is cancelled by a compensating term arising from the integration of the Faddeev–Popov determinant that arises in the path-integral measure when gauge fixing Dasgupta:2001ue ; Ambjorn:2002gr . A cancellation mechanism similar to this should apply also in nonlocal gravity, since it amounts to a nonperturbative upgrade of the argument exposed in section 4 based on the same physical principle of gauge invariance. This aspect will deserve a check in the future.

### v2

@@ 3 Tree-level graviton propagator
can be computed in a standard albeit lengthy way. The expansion at second order of the action (4) yields the Lagrangian [^2] Here is how to compare the conventions and results of Accioly:2002tz with ours. Spacetime signature is mostly minus in Accioly:2002tz and mostly plus for us. Also, their Ricci scalar $R$ is $-R$ for us, so that there is an extra overall -- sign in the total action and in its coefficients $\alpha=-M_{\text{\tiny Pl}}^{2}\gamma_{0}$, $\beta=-M_{\text{\tiny Pl}}^{2}\gamma_{2}$ and $\gamma=-M_{\text{\tiny Pl}}^{2}\gamma_{4}$. Then, $b=\beta\kappa^{2}/2=-2\gamma_{2}$, $c=\alpha/\beta=\gamma_{0}/\gamma_{2}$ and $d=\gamma\kappa^{2}/2=-2\gamma_{4}$, where $\kappa=2/M_{\text{\tiny Pl}}$. Finally, the graviton in Accioly:2002tz is dimensionful and defined as $g_{\mu\nu}=\eta_{\mu\nu}+\kappa h_{\mu\nu}$, while our $h_{\mu\nu}$ is dimensionless. Therefore, our coefficients $c_{i}\in\{c_{1},c_{2},c_{0},\bar{c}_{0},\bar{\bar{c}}_{0}\}$ are related to the coefficients $x_{i}\in\{x_{1},x_{2},x_{0},\bar{x}_{0},\bar{\bar{x}}_{0}\}$ in Accioly:2002tz by $c_{i}(k^{2})=-(M_{\text{\tiny Pl}}^{2}/4)\,x_{i}(k^{2})$. The expressions in Accioly:2002tz immediately generalize to the weakly nonlocal case with non-trivial form factors and we have recalculated them anew.
@@
Contrary to $\mathcal{O}^{\rm K}$, the kinetic operator $\mathcal{O}=\mathcal{O}^{\rm K}+\mathcal{O}^{\rm gf}$ is invertible. In the basis of Barnes–Rivers projectors Riv64 ; Bar65 ; VanNieuwenhuizen:1973fi and in momentum space, it is
@@
This expression is equivalent to ${\mathcal{O}}_{\rm K}^{-1}={P}^{(2)}/c_{2}+{P}^{(0)}/c_{0}$ when setting $\lambda_{1}=\lambda_{2}=\lambda_{3}=0$. In general, different gauge choices become handy when exploring different aspects of the theory, for instance, when showing that the spin-0 mode with the wrong sign does not propagate or when checking positivity of the eigenvalues of the kinetic operator in Euclidean signature. Some common gauge choices are the Julve–Tonin gauge ($\lambda=0$) Julve:1978xn , the de Donder gauge ($\lambda=1/2$, $\lambda_{2}=0=\lambda_{3}$) and the Feynman gauge ($\lambda=1/2$, $\lambda_{1}=1$, $\lambda_{2}=0=\lambda_{3}$) Accioly:2002tz .
@@ 4 Conformal instability
where the fields $\Phi_{i}$ include Euclidean metric $g_{\mu\nu}^{\rm E}$ and $S_{\rm E}$ is the Euclidean action. This theory can suffer from a conformal instability problem Hawking:1976ja ; Gibbons:1978ac ; Ham09 , which we illustrate in Einstein gravity. A conformal transformation of the metric
@@
in the Ricci term $R$ produces a kinetic term for the conformal factor $\Omega$ with the wrong sign. The Lorentzian action is
$S=\frac{M_{\text{\tiny Pl}}^{2}}{2}\int d^{D}x\,\sqrt{|g|}\,R=\frac{M_{\text{\tiny Pl}}^{2}}{2}\int d^{D}x\,\sqrt{|\hat{g}|}\,\Omega^{D-4}\left[\Omega^{2}\hat{R}+(D-1)(D-2)\hat{\partial}_{\mu}\Omega\hat{\partial}^{\mu}\Omega\right],$ (69)
@@
Therefore, the Euclidean path integrand $\exp(iS)=\exp(-S_{\rm E})$ has an anti-Gaussian term $\sim\exp[\int(\partial\omega)^{2}]$ and is unbounded from above. To avoid this arbitrarily large instability, one can deform the integration contour over the fields, in particular, taking a functional measure $[\mathcal{D}g_{\mu\nu}]$ over complex Euclidean metrics (complex conformal factor), in such a way as to get a convergent result Gibbons:1978ac ; Gibbons:1978ji ; Christensen:1979iy ; tHooft:2010xlr ; tHooft:2010mvw ; Marolf:2022ntb .
@@
The physical significance of this procedure was criticized in Schleich:1987fm ; Hartle:2020glw ; Mazur:1989by and motivated another approach, where the fundamental definition of the path integral is Lorentzian and one derives (instead of assuming) the correct Euclidean path integral by integrating over a suitable contour in field space with real Lorentzian metrics Schleich:1987fm ; Hartle:2020glw ; Mazur:1989by ; Marolf:2022ybi . Nonlocal quantum gravity is defined with a Lorentzian path integral, so that the issue of the conformal instability can be solved in this way. At least, there is no reason to suspect that this resolution, successful in other theories, should not work also here. As is clear from the construction at the beginning of section 2 and in section 2.1, all dynamics-independent tools developed to solve the conformal instability problem are available also in nonlocal quantum gravity. In particular, the York decomposition York:1973ia ; York:1974psa of metric perturbations $h_{\mu\nu}$ into transverse-traceless, scalar and gauge (coordinate) modes holds, and we also have a DeWitt metric defining the quadratic form $\langle h_{\mu\nu},h_{\sigma\tau}\rangle$ DeWitt:1967yk ; DeWitt:1967ub ; DeWitt:1967uc . These tools allow one to handle the Jacobian factors in the path-integral measure as usual and to factorize them into Gaussian integrals over each of the metric York components. Then, since entire nonlocal form factors do not break Lorentz and diffeomorphism invariance nor hide any extra pole in field redefinitions, all the manipulations of Mazur:1989by and of its non-perturbative extension Dasgupta:2001ue leading to a well-defined continuation to a Euclidean path integral can be applied to the nonlocal action.
@@
In this paper, we will not discuss the resolution of the conformal instability problem at the non-perturbative level until section 5 because we mainly focus on perturbation theory, where the Lorentzian path integral is employed. As described in section 2.5, amplitudes are calculated in Euclidean momenta and then analytically continued to imaginary Euclidean (i.e., real Lorentzian) external energies. However, at the non-perturbative level, the Lorentzian and the Euclidean theory are distinguishable because the path integral, with its functional measure and the space of metrics on which one integrates Marolf:2022ybi ; Dong:2019piw , differs in the Euclidean and Lorentzian formulations. Redefining and analytically continuing the Lorentzian momenta of the fields appearing in the functional measure of the Lorentzian path integral produces a measure which is functionally inequivalent to the one obtained by analytically continuing the Euclidean momenta of the fields appearing in the measure of the Euclidean path integral. To put it very schematically,
Lorentzian path integral $\coloneqq$ $\int_{\textrm{metrics}\,\,(-,+,\dots,+)}\mathcal{D}g_{\mu\nu}\,e^{iS}$
with Efimov prescription on $k^{0}$ integration
$=$ $\int_{\textrm{metrics}\,\,(-,+,\dots,+)}\mathcal{D}g_{\mu\nu}\,e^{-S_{\rm E}}$
with Efimov prescription on $k_{D}=-ik^{0}$ integration
$=$ $\int_{\textrm{metrics}\,\,(+,+,\dots,+)}\mathcal{D}g_{\mu\nu}^{\rm E}\,\mathcal{J}\,e^{-S_{\rm E}}$
with Efimov prescription on $k_{D}$ integration
$\neq$ $\int_{\textrm{metrics}\,\,(+,+,\dots,+)}\mathcal{D}g_{\mu\nu}^{\rm E}\,e^{-S_{\rm E}}$
with naive analytic continuation
[…]
where the action is expressed as an integral in momentum space and the Jacobian $\mathcal{J}=\mathcal{D}g_{\mu\nu}/\mathcal{D}g_{\sigma\tau}^{\rm E}$ can be inferred from Dasgupta:2001ue . In particular, nonlocal quantum gravity is defined by the left-hand side of the first line of (LABEL:scheme), not by the last line.
@@
From this discussion, it stems that the conformal instability problem does not affect perturbative results because, in general, the Lorentzian perturbation theory is well defined Ham09 ; Ordonez:1985kz . Let us expand on this statement. In order for the Lorentzian path integral to be well defined in perturbation theory, the tree-level graviton kinetic term in Feynman prescription $k^{2}\to k^{2}-i\epsilon$ must generate a Gaussian term in the path integral that makes it convergent. If this was the case, then we would have convergence order by order in perturbation theory because, at higher orders, one only has to include vertex insertions, which do not spoil the tree-level convergence property. This is because all correlation functions in perturbation theory are written as an expansion of the interactions $S_{\rm int}$: schematically for a generic field $\Phi$,
@@
where $S_{0}=\int\Phi\mathcal{O}\Phi$ and $\mathcal{O}=\mathcal{O}^{\rm K}+\mathcal{O}^{\rm gf}$ is the tree-level kinetic term (inverse of the propagator), where $\mathcal{O}^{\rm K}$ is the kinetic term coming from the bare action and $\mathcal{O}^{\rm gf}$ is a contribution coming from the gauge-fixing action. In order to explicitly compute any perturbative amplitude, we should commute the functional integral with the sum. This issue is related to the Borel summability of the perturbative series that deserves to be investigated also in gravity and, in particular, in a finite theory of quantum gravity in the QFT framework.
If one naively considered only the gauge-invariant part $\mathcal{O}^{-1}_{\rm K}$ of the propagator, the convergence problem would persist after the $k^{2}\to k^{2}-i\epsilon$ prescription because the spin-2 and spin-0 modes in $\mathcal{O}^{\rm K}$ have opposite sign, so that the spin-0 mode would generate an anti-Gaussian profile. Thus, the role of $\mathcal{O}^{\rm gf}$ cannot be ignored. If, after a Weyl transformation (68), a suitable gauge choice exists such that the eigenvalues of the kinetic operator of the conformal factor $\Omega$ are all zero or negative in Euclidean signature, then there is convergence of the Euclidean path integral at the perturbative level.
To show this, it is sufficient to consider the gravitational sector of nonlocal quantum gravity and the generic action (4) where now $\gamma_{0}$, $\gamma_{2}$ and $\gamma_{4}$ are generic form factors which vanish for Einstein gravity and are constant in Stelle gravity, and $S_{\rm gf}$ is the gauge-fixing action. In a Weyl invariant theory, the metric $g_{\mu\nu}$ composing the measure and the curvature tensors is the one before making explicit the dilaton dependence. Decompose the metric as in (57) and derive the perturbed action for the dimensionless spin-2 field $h_{\mu\nu}$ (the graviton polarization modes are inside this object). On Minkowski background, the kinetic term is $\propto h^{\mu\nu}\mathcal{O}_{\mu\nu\sigma\tau}h^{\sigma\tau}$, where the operator $\mathcal{O}_{\mu\nu\sigma\tau}$ is given in section 3 and is a generalization of the higher-derivative expression found in Accioly:2002tz .
@@
where we used eq. (60), the Barnes–Rivers projectors (61) and the ensuing properties
@@
Intuitively, only the spin-0 projectors survive because they act on a scalar. From the expressions (62) for the coefficients $c_{0}$, $\bar{c}_{0}$ and $\bar{\bar{c}}_{0}$, in momentum space we have
@@
If we ignored the gauge-fixing terms and set $\lambda=\lambda_{1}=\lambda_{2}=0$, then $\mathcal{K}>0$ in Euclidean signature for non-negative form factors $\gamma_{0,2}^{\prime}\geqslant 0$ and the Euclidean action $S_{\rm E}\sim-\int\phi\mathcal{K}\phi$ would be unbounded from above. This is the conformal instability problem. However, by virtue of gauge invariance, if we can find a range of gauge choices where $\mathcal{K}\leqslant 0$ for all momenta $k$, then we can conclude that such problem is a gauge artefact and that the scalar mode $\phi$ does not propagate ($\mathcal{K}=0$) or that it propagates with a kinetic term of the “straight” sign ($\mathcal{K}<0$). In general, for $\gamma_{2}^{\prime}\neq 0$ this happens when
@@
The first of these conditions appears in all theories with an Einstein–Hilbert term and implies that both $\lambda$ and $\lambda_{1}$ can be chosen to be constant. In $D=4$ dimensions, $\lambda_{1}\geqslant 3/(4\lambda-1)^{2}$ and infinitely many gauges with $\lambda\neq 1/4$ can fulfill eq. (77), including Julve–Tonin ($\lambda=0$) and de Donder ($\lambda=1/2$, $\lambda_{2}=0=\lambda_{3}$), but not the Feynman gauge ($\lambda=1/2$, $\lambda_{1}=1$, $\lambda_{2}=0=\lambda_{3}$).
@@ 5 Discussion
To date, and modulo some very preliminary exceptions Ghoshal:2022mnj , nonlocal quantum gravity has been studied only with perturbative techniques in relation with scattering amplitudes, perturbative unitarity, renormalization, black-hole solutions and cosmology. However, the conformal instability problem could arise when considering non-perturbative processes and the full path integral is not expanded in terms of the interactions. As in the perturbative case, once the Euclidean path integral is derived from the Lorentzian one, a non-trivial choice of the gravitational measure of the path integral can make it convergent. The physical interpretation is that this measure would enhance the weight of strong metric fluctuations pushing away physical fluctuations from the conformal unbounded abyss Ham09 . This seems indeed to be the case, both on a lattice at strong coupling Hamber:1984kx ; Berg:1985ka and from a non-perturbative extension Dasgupta:2001ue ; Ambjorn:2002gr of the perturbative calculation of Mazur:1989by . In the case of the Einstein–Hilbert action, after a Wick rotation of the non-perturbative Lorentzian path integral defined on a causal dynamical triangulation (CDT), it turns out that the term responsible for the conformal instability is cancelled by a compensating term arising from the integration of the Faddeev–Popov determinant that arises in the path-integral measure when gauge fixing Dasgupta:2001ue ; Ambjorn:2002gr . A cancellation mechanism virtually identical to this should apply also in nonlocal gravity, since it amounts to a non-perturbative upgrade of the argument exposed in section 4 based on the same physical principle of gauge invariance. In particular, the calculation of the measure and the cancellation mechanism in (Dasgupta:2001ue, , section 4) apply verbatim also to nonlocal quantum gravity because the conformal divergence stems from the $R$ term in the action, eqs. (69) and (70); in fact, a conformal transformation of $O(R^{2},R_{\mu\nu}R^{\mu\nu})$ operators with or without nonlocal form factors does not produce terms with the opposite sign. Thus, keeping the leading conformal divergence from the Einstein–Hilbert term $R$ and integrating over the conformal mode after a field redefinition introduced in Distler:1988jt , the compensation of the unbounded term by the functional determinant takes place. Numerical analyses in CDT support this conclusion for Einstein gravity. Upon discretization, the action is bounded from below for a fixed discrete spacetime volume. The minimum of the action happens when the ratio $N_{22}/N_{3}$ between the number $N_{22}$ of so-called 2-2 tetrahedra and the total number $N_{3}$ of tetrahedra is minimal. Numerical simulations in three dimensions show that the expectation value $\langle N_{22}/N_{3}\rangle$ stays positive and away from zero in the continuum limit Ambjorn:2000dja ; Ambjorn:2001br , meaning that the kinetic term of the conformal mode never dominates the dynamics and its contribution to the path integral is suppressed in the continuum Lorentzian limit Dasgupta:2001ue ; Ambjorn:2002gr . All of this is relevant also for nonlocal quantum gravity, since CDT is not an independent proposal of quantum gravity but, rather, a regularization method that can be applied to any path integral with a diffeomorphism-invariant action. Entire form factors do not add any extra poles nor change the sign of kinetic terms and asymptotically local quantum gravity on a CDT should inherit the same qualitative properties found for Einstein gravity. These theoretical and numerical aspects will deserve a check in the future in the nonlocal case.